An alloy of copper and tin is 20% copper. How many pounds of copper must be added to 20 pounds of the alloy to get an alloy that is 50% copper?
Answer by ankor@dixie-net.com(22740)

An alloy of copper and tin is 20% copper.
How many pounds of copper must be added to 20 pounds of the alloy to get an alloy that is 50% copper?
:
let x = amt of pure copper required
.20(20) + x = .50(x+20)
4 + x = .5x + 10
x - .5x = 10 - 4
.5x = 6
x = 12 lb of pure copper to accomplish this
;
;
Check:
.2(20) + 12 = .5(12+20))
4 + 12 = 16